2) (3 pts) Use mathematical induction to show that when n is an exact power of 2, the solution of the recurrence 2, ifn=2 T(n) =127G)+n, ifn=2.for k > 1 ISI(72) = n lg n.
Answer #1
T(n) = 2T(n/2) + n
T(n) = 2(2T(n/4) + n/2) + n
T(n) = 4T(n/4) + n + n
T(n) = 8T(n/8) + n + n + n
.....
.....
.....
T(n) = nT(n/n) + n + ......... + n + n + n
T(n) = n + n + ......... + n + n + n [log(n) times]
T(n) = n log(n)
T(n) = O(n log(n))